# Matchings in Matroids over Abelian Groups, III

### v1

@@ Abstract.
In an abelian group $G$, a matching is a bijection $f\colon A\to B$ between finite subsets $A,B\subseteq G$ such that $a+f(a)\notin A$ for all $a\in A$. We say that $G$ has the matching property if every pair of finite subsets $A,B\subseteq G$ with $|A|=|B|$ and $0\notin B$ admits such a matching. This paper develops matroidal analogues of classical results on group matchings. By embedding matroid ground sets in $G$, we introduce base matchings between matroid bases, recovering the group-theoretic setting in the uniform case, and derive structural and combinatorial criteria for their existence. Our methods blend techniques from matroid theory, group theory, and additive number theory. Our main focus is on paving matroids, a class conjectured to constitute asymptotically almost all matroids. We prove symmetric self-matchability for all paving matroids, extend asymmetric results via the hyperplane-nullity parameter, and connect stressed hyperplanes to matchability through relaxation, bridging paving and uniform matroids. This paper continues a line of research initiated in [4, 3], yet is written to be self-contained and may be read independently.
Key words and phrases:
critical pairs, matchable bases, paving matroids, progressions, stressed hyperplanes
2020 Mathematics Subject Classification:
Primary: 05B35; Secondary: 05D15, 11B13.
@@ 1.1. Overview
A geometric model for a class of bipartite graphs was introduced in [9], where a particular type of perfect matchings, termed an acyclic matching, was defined and shown to exist for a subclass of these graphs using geometric arguments. The existence of such acyclic matchings implies the non-vanishing of the determinant of certain weighted biadjacency matrices. This matching framework was applied to a question posed by E. K. Wakeford in 1920 [23], which concerns determining the sets of monomials that can be eliminated from a generic homogeneous polynomial via linear transformations of the variables. In a special case involving the canonical form of symmetric tensors, Fan and Losonczy reduced Wakeford’s problem to the existence of acyclic matchings in $\mathbb{Z}^{n}$. This property was established for $\mathbb{Z}^{n}$ by Alon, Fan, Kleitman, and Losonczy [5], and subsequently characterized completely for all abelian groups in [2].
@@
Matchings in abelian groups: Throughout, we use the notation $n]:=\{1,\dots,n\}$ for any positive integer $n$. We also write $(G,+)$ for an abelian group with neutral element 0. Let $p(G)$ represent the smallest cardinality of a nontrivial (i.e., non-zero) finite subgroup of $G$. If $G$ has no nontrivial finite subgroup, we make the convention $p(G)=\infty$. Let $A$ and $B$ be finite subsets of $G$ with equal cardinality, such that $0\notin B$. A matching from $A$ to $B$ is a bijection $f:A\to B$ satisfying the condition that $a+f(a)\notin A$ for every $a\in A$. The conditions $|A|=|B|$ and $0\notin B$ are necessary for the existence of such a bijection. A matching is called symmetric if $A=B$, and asymmetric otherwise. If there exists a matching from $A$ to $B$, then $A$ is said to be matched or matchable to $B$. An abelian group $G$ is said to satisfy the matching property if for every pair of finite subsets $A,B\subseteq G$ with $|A|=|B|$ and $0\notin B$, the set $A$ is matched to $B$.
@@
A connection to matchings in graphs: While there are meaningful connections between the notion of matchings explored in this paper and the classical matching theory in graphs introduced by Philip Hall in 1935, the two concepts are fundamentally distinct. In particular, to address a linear algebra problem posed by Wakeford [23], a bipartite graph $\mathcal{G}=(V(\mathcal{G}),E(\mathcal{G}))$ was constructed in [9] using two sets $A\subseteq\mathbb{Z}^{n}$ and $B\subseteq\mathbb{Z}^{n}$, as follows:
Since $A$ and $B$ may not be disjoint, each element $a\in A$ is associated with a symbol $x_{a}$, and each $b\in B$ with a symbol $y_{b}$. Define $X=\{x_{a}\mid a\in A\}$ and $Y=\{y_{b}\mid b\in B\}$. The bipartite graph $\mathcal{G}$ has vertex set $V(\mathcal{G})=X\cup Y$, and an edge connects $x_{a}\in X$ to $y_{b}\in Y$ if and only if $a+b\notin A$.
@@ 1.2. Matching in matroids
The rank of a matroid $M$, denoted by $r(M)$, is defined to be $r_{M}(E(M))$. A set $X\subseteq E$ is called dependent if it is not independent. A maximal independent set is called a basis. It follows from the augmentation property that every two bases have the same size. A minimal dependent set is called a circuit. An element that forms a single-element circuit of $M$ is called a loop. A matroid is called loopless if it does not have a loop. If $X\subseteq E$, the closure of $X$ is denoted by ${\it{cl}}(X)$ and defined as follows:
${\it{cl}}(X)=\{x\in E:r(X)=r(X\cup\{x\})\}.$
The closure operator satisfies the following properties:
  • For all subset $X$ of $E$, $X\subseteq{\it{cl}}(X)$.
  • If r(X)=r(M)1r(X)=r(M)-1, then ${\it{cl}}(X)$ is a hyperplane of $M$.
@@
    We say that $M$ is matched to $N$ if for every basis $\mathcal{M}$ of $M$ there exists a basis $\mathcal{N}$ of $N$, so that $\mathcal{M}$ is matched to $\mathcal{N}$.
Interplay between matchings in groups and matchings in matroids
  • Matchability in the group setting can be viewed as a special case of matchability in the matroid setting. Let $A$ and $B$ be two finite nonempty subsets of an abelian group $G$, each of cardinality $n$, and assume $0\notin B$. Define a uniform matroid $M\cong U_{n,n}$ on the ground set $A$, and similarly define $N\cong U_{n,n}$ on the ground set $B$. Then $A$ is matched to $B$ in the group-theoretic sense if and only if $M$ is matched to $N$ in the matroid-theoretic sense. Note that uniform matroids form a very narrow subclass of all matroids, and $U_{n,n}$ represents a particularly special case.
  • Adopting the notation from Definition 1.4, suppose $\mathcal{M}$ is matched to $\mathcal{N}$ as matroid bases. Then for each $1\leq i\leq n$, we have $a_{i}+b_{i}\notin E(M)$, and in particular $a_{i}+b_{i}\notin\mathcal{M}$. This implies that the map $a_{i}\mapsto b_{i}$ defines a matching in the group-theoretic sense between the subsets $\mathcal{M}$ and $\mathcal{N}$ of $G$. In this way, the definition of matchings in matroids is compatible with the classical notion of matchings in abelian groups.
Compatibility of matchings in matroids with the vector space setting
The definition in Definition 1.4 is also consistent with the notion of matchability in vector spaces over field extensions, as introduced by Eliahou and Lecouvey [8]. Let $K\subseteq F$ be a field extension, and let $A,B\subseteq F$ be $n$-dimensional $K$-vector subspaces. Suppose $\mathcal{A}=\{a_{1},\dots,a_{n}\}$ and $\mathcal{B}=\{b_{1},\dots,b_{n}\}$ are ordered bases of $A$ and $B$, respectively. Then $\mathcal{A}$ is said to be matched to $\mathcal{B}$ if
@@ 2.1. Notions and results on paving matroids
A $d$-partition of a finite set $E$ is a collection $\mathcal{S}$ of subsets of $E$, all of which of size at least $d$, so that the intersection of any two sets in $\mathcal{S}$ is of size at most d1d-1. The set $\mathcal{S}=\{E\}$ is a $d$-partition for any $d$, which we call a trivial $d$-partition. The connection between paving matroids and d-partitions is given by the following result whose proof can be found in [18].
@@
[18] If $M$ is a paving matroid of rank $n\geq 2$, then its hyperplanes form a non-trivial (n1)(n-1)-partition of $E(M)$. That is, the intersection of any two hyperplanes in $M$ is of size at most n2n-2.
@@
Conversely, assume that ${\it null}(\mathcal{H}_{M})\leq 1$. Let $A$ be an $n$-subset of $E(M)$ that is not a basis. We claim that $A$ is a circuit-hyperplane. Since $M$ is paving then $A$ is a circuit with r(A)=n1r(A)=n-1 Consider the hyperplane $H=\operatorname{cl}(A)$. Then ${\it null}(H)\leq 1$. We have
@@
Hence all inequalities are equalities, which forces $A=H$. Thus $A$ is a hyperplane. Since $A$ is a circuit-hyperplane, by Theorem 2.2, $M$ is sparse paving. ∎
@@ 2.2. Abelian additive theory
We start with a result of Kneser[17, Theorem 4.3] providing a lower bound on the size of $A+B$.
@@
Let $G$ be an abelian group, $x,a\in G$ and let $k$ be a positive integer. A progression of length $k$ with difference $x$ and initial term $a$ is a subset of $G$ of the form $\{a,a+x,a+2x,...,a+(k-1)x\}.$. We say $A$ is a semi-progression if $A\setminus\{a\}$ is a progression, for some $a\in A$. In the following lemma due to Kemperman, a certain family of critical pairs is characterized. Note that a pair $(A,B)$ of finite subsets of a group $G$ is called critical if |G|>|A+B|=|A|+|B|1|G|>|A+B|=|A|+|B|-1.
@@ 3.1. Matching paving matroids over abelian groups (symmetric case)
Note that, for any pair of matroids $M,N$, the conditions $0\notin E(N)$ and $r(M)=r(N)$ are necessary for $M$ to be matched to $N$ (see [4, Proposition 1.7]). Thus, the “only if” implication in Theorem 3.2 is immediate. We now establish the reverse implication.
@@
a contradiction. This completes the proof of the claim. ∎
@@
We now show that $M$ is matched to itself by demonstrating that each ordered basis of $M$ can be matched to another ordered basis of $M$:
  • $\{1,2,5\}$ is matched to $\{5,4,1\}$,
  • $\{1,3,5\}$ is matched to $\{5,4,1\}$,
  • $\{1,4,5\}$ is matched to $\{5,4,1\}$,
  • $\{2,3,5\}$ is matched to $\{5,4,1\}$,
  • $\{2,4,5\}$ is matched to $\{5,4,1\}$,
  • $\{3,4,5\}$ is matched to $\{5,4,1\}$.

### v2

@@ Abstract.
This paper develops matroidal analogues of classical results on matchings in abelian groups. By embedding matroid ground sets in an abelian group, we introduce base matchings between matroid bases, recover the group-theoretic setting in the uniform matroid case, and derive structural and combinatorial criteria for their existence. Our main focus is on paving matroids. We prove self-matchability for paving matroids, extend asymmetric matchability results using the hyperplane-nullity parameter, and show that stressed hyperplanes provide a natural route to matchability through relaxation.
@@ 1.1. Overview
In an abelian group $G$, a matching is a bijection $f\colon A\to B$ between finite subsets $A,B\subseteq G$ such that $a+f(a)\notin A$ for all $a\in A$. We say that $G$ has the matching property if, for every pair of finite subsets $A,B\subseteq G$ with $|A|=|B|$ and $0\notin B$, there exists such a matching. A geometric model for a class of bipartite graphs was introduced in [11]. In that work, a special kind of perfect matching, called an acyclic matching, was defined and shown to exist for a subclass of these graphs by geometric methods. The existence of acyclic matchings implies the nonvanishing of the determinant of certain weighted biadjacency matrices. This matching framework was later applied to a question posed by E. K. Wakeford in 1920 [25], concerning which sets of monomials can be eliminated from a generic homogeneous polynomial by linear changes of variables. In a special case related to canonical forms of symmetric tensors, Fan and Losonczy reduced Wakeford’s problem to the existence of acyclic matchings in $\mathbb{Z}^{n}$. This property was established for $\mathbb{Z}^{n}$ by Alon, Fan, Kleitman, and Losonczy [7], further developed in [1, 4], and later completely characterized for all abelian groups in [3].
@@
Matchings in abelian groups. Throughout, we use the notation $n]:=\{1,\dots,n\}$ for any positive integer $n$. We also write $(G,+)$ for an abelian group with neutral element 0. Let $p(G)$ represent the smallest cardinality of a nontrivial (i.e., non-zero) finite subgroup of $G$. If $G$ has no nontrivial finite subgroup, we make the convention $p(G)=\infty$. Let $A$ and $B$ be finite subsets of $G$ with equal cardinality, such that $0\notin B$. A matching from $A$ to $B$ is a bijection $f:A\to B$ satisfying the condition that $a+f(a)\notin A$ for every $a\in A$. The conditions $|A|=|B|$ and $0\notin B$ are necessary for the existence of such a bijection. A matching is called symmetric if $A=B$, and asymmetric otherwise. If there exists a matching from $A$ to $B$, then $A$ is said to be matched or matchable to $B$. An abelian group $G$ is said to satisfy the matching property if for every pair of finite subsets $A,B\subseteq G$ with $|A|=|B|$ and $0\notin B$, the set $A$ is matched to $B$.
@@
A connection to matchings in graphs. While there are meaningful connections between the notion of matchings explored in this paper and the classical matching theory in graphs introduced by Philip Hall in 1935, the two concepts are fundamentally distinct. In particular, in order to address a linear algebra problem posed by Wakeford [25], Fan [11] constructed a bipartite graph
$\mathcal{G}=(V(\mathcal{G}),E(\mathcal{G}))$
from two sets $A,B\subseteq\mathbb{Z}^{n}$. To each element $a\in A$ one associates a symbol $x_{a}$, and to each element $b\in B$ a symbol $y_{b}$. Let
$X=\{x_{a}\mid a\in A\}\qquad\text{and}\qquad Y=\{y_{b}\mid b\in B\}.$
The graph $\mathcal{G}$ is then defined to have vertex set
$V(\mathcal{G})=X\cup Y,$
where an edge joins $x_{a}\in X$ to $y_{b}\in Y$ if and only if
$a+b\notin A.$
@@ 1.2. Matching in Matroids
The rank of a matroid $M$, denoted by $r(M)$, is defined to be $r_{M}(E(M))$. A set $X\subseteq E$ is called dependent if it is not independent. A maximal independent set is called a basis. It follows from the augmentation property that every two bases have the same size. A minimal dependent set is called a circuit. An element that forms a single-element circuit of $M$ is called a loop. A matroid is called loopless if it does not have a loop.
@@
If $X\subseteq E$, the closure of $X$ is denoted by ${\it{cl}}(X)$ and defined as follows:
${\it{cl}}(X)=\{x\in E:r(X)=r(X\cup\{x\})\}.$
The closure operator satisfies the following properties:
  • For all subsets $X$ of $E$, $X\subseteq{\it{cl}}(X)$.
  • If r(X)=r(M)1r(X)=r(M)-1, then ${\it{cl}}(X)$ is a hyperplane of $M$.
@@
    We say that $M$ is matched to $N$ if for every basis $\mathcal{M}$ of $M$ there exists a basis $\mathcal{N}$ of $N$ such that $\mathcal{M}$ is matched to $\mathcal{N}$.
Note that adopting the notation from Definition 1.4, suppose $\mathcal{M}$ is matched to $\mathcal{N}$ as matroid bases. Then for each $1\leq i\leq n$, we have $a_{i}+b_{i}\notin E(M)$, and in particular $a_{i}+b_{i}\notin\mathcal{M}$. This implies that the map $a_{i}\mapsto b_{i}$ defines a matching in the group-theoretic sense between the subsets $\mathcal{M}$ and $\mathcal{N}$ of $G$. In this way, the definition of matchings in matroids is compatible with the classical notion of matchings in abelian groups.
Compatibility of matchings in matroids with the vector space setting. Definition 1.4 is consistent with the notion of matchability in vector spaces over field extensions, as introduced by Eliahou and Lecouvey [10]. Let $K\subseteq F$ be a field extension, and let $A,B\subseteq F$ be $n$-dimensional $K$-vector subspaces. Suppose $\mathcal{A}=\{a_{1},\dots,a_{n}\}$ and $\mathcal{B}=\{b_{1},\dots,b_{n}\}$ are ordered bases of $A$ and $B$, respectively. Then $\mathcal{A}$ is said to be matched to $\mathcal{B}$ if
@@
The point of passing from sparse paving to paving matroids is not merely to enlarge the class of matroids for which matching results are known. In sparse paving matroids (with rank $n$), the only dependent $n$-sets are circuit-hyperplanes, so the obstruction is essentially binary. In other words, the failure of an $n$-subset to be a basis is controlled by a single type of obstruction, namely circuit-hyperplanes. By contrast, in paving matroids, hyperplanes may have larger nullity, and our results show that matchability is controlled by this richer geometric parameter. In this way, the paper shows that the key mechanism governing matroid matchings is the structure of hyperplanes, rather than near-uniformity alone.
@@ 2.1. Notions and Results on Paving Matroids
A $d$-partition of a finite set $E$ is a collection $\mathcal{S}$ of subsets of $E$, all of which are of size at least $d$, so that the intersection of any two distinct sets in $\mathcal{S}$ is of size at most d1d-1. The set $\mathcal{S}=\{E\}$ is a $d$-partition for any $d$, which we call a trivial $d$-partition. The connection between paving matroids and d-partitions is given by the following result, whose proof can be found in [20].
@@
[20] If $M$ is a paving matroid of rank $n\geq 2$, then its hyperplanes form a non-trivial (n1)(n-1)-partition of $E(M)$. That is, the intersection of any two distinct hyperplanes in $M$ is of size at most n2n-2.
@@
Conversely, assume that ${\it null}(\mathcal{H}_{M})\leq 1$. Let $A$ be an $n$-subset of $E(M)$ that is not a basis. We claim that $A$ is a circuit-hyperplane. Since $M$ is paving, then $A$ is a circuit with r(A)=n1r(A)=n-1. Consider the hyperplane $H=\operatorname{cl}(A)$. Then ${\it null}(H)\leq 1$. We have
@@
Hence, all inequalities are equalities, which forces $A=H$. Thus $A$ is a hyperplane. Since $A$ is a circuit-hyperplane, by Theorem 2.2, $M$ is sparse paving. ∎
@@ 2.2. Abelian Additive Theory
We start with a result of Kneser[19, Theorem 4.3] providing a lower bound on the size of $A+B$.
@@
Let $G$ be an abelian group, $x,a\in G$ and let $k$ be a positive integer. A progression of length $k$ with difference $x$ and initial term $a$ is a subset of $G$ of the form $\{a,a+x,a+2x,...,a+(k-1)x\}$. We say $A$ is a semi-progression if $A\setminus\{a\}$ is a progression, for some $a\in A$. In the following lemma due to Kemperman, a certain family of critical pairs is characterized. Note that a pair $(A,B)$ of finite subsets of a group $G$ is called critical if |G|>|A+B|=|A|+|B|1|G|>|A+B|=|A|+|B|-1.
@@ 3. Main Results
Note that matchability in the group setting can be viewed as a special case of matchability in the matroid setting. Let $A$ and $B$ be two finite nonempty subsets of an abelian group $G$, each of cardinality $n$, and assume $0\notin B$. Define a uniform matroid $M\cong U_{n,n}$ on the ground set $A$, and similarly define $N\cong U_{n,n}$ on the ground set $B$. Then $A$ is matched to $B$ in the group-theoretic sense if and only if $M$ is matched to $N$ in the matroid-theoretic sense. Note that uniform matroids form a very narrow subclass of all matroids, and $U_{n,n}$ represents a particularly special case.
The importance of paving matroids is not only that they extend sparse paving matroids, but that they make the structural source of matchability more visible. Since the classical group-theoretic notion of matching is recovered in the uniform case, paving matroid matchability extends the original problem of matchings within groups. Passing to all paving matroids, therefore, helps separate features of matchability that depend on near-uniformity from those that persist in a richer dependence structure. Our results show that, in paving matroids, matchability is governed by geometric features such as hyperplane intersections, hyperplane nullity, and stressed hyperplanes. This both explains why the theory extends beyond the sparse paving case and places the classical group case within a broader structural framework. Since paving matroids can also be related back to uniform matroids through relaxation (Theorem 3.23), the group-theoretic setting appears not as an isolated phenomenon, but as the first instance of a more general theory of matchings.
@@ 3.1. Matching Paving Matroids over Abelian Groups (Symmetric Case)
Note that, for any pair of matroids $M,N$, the conditions $0\notin E(N)$ and $r(M)=r(N)$ are necessary for $M$ to be matched to $N$ (see [6, Proposition 1.7]). Thus, the “only if” implication in Theorem 3.2 is immediate. We now establish the reverse implication.
@@
which is a contradiction. This completes the proof of the claim. ∎
@@
Remark 3.6.
The significance of Theorem 3.2 is that it shows the classical self-matching criterion from abelian groups (Theorem 1.1) is not confined to sets or to uniform matroids. Even in a paving matroid, where substantial dependence may occur, the only obstruction to symmetric self-matchability remains the trivial additive obstruction $0\in E(M)$. Thus, the theorem identifies the paving condition as a structural setting in which matroidal dependence does not create new symmetric matching obstructions. In particular, it shows that the source of self-matchability lies not merely in near-uniformity, but in the hyperplane geometry of paving matroids.
@@
We use the construction of Theorem 3.2 to show that $M$ is matched to itself. First, consider the matching
[…]
in the group setting.
  • Consider the basis $\mathcal{M}=\{1,2,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{5,4,1\},$
    which is a basis of $M$, and $\mathcal{M}$ is matched to $\mathcal{N}$.
  • Consider the basis $\mathcal{M}=\{1,3,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{5,3,1\},$
    which is a basis of $M$, and $\mathcal{M}$ is matched to $\mathcal{N}$.
  • Consider the basis $\mathcal{M}=\{1,4,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{5,2,1\},$
    which is a basis of $M$, and $\mathcal{M}$ is matched to $\mathcal{N}$.
  • Consider the basis $\mathcal{M}=\{2,3,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{4,3,1\},$
    which is not a basis of $M$. Choose $a_{3}=5$ and $x=5$ as in Claim 3.3, and form the basis
    $\mathcal{P}=\{4,3,5\}$
    of $M$. Then $\mathcal{M}$ is matched to $\mathcal{P}$.
  • Consider the basis $\mathcal{M}=\{2,4,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{4,2,1\},$
    which is not a basis of $M$. Choose $a_{1}=2$ and $x=5$ as in Claim 3.3, and form the basis
    $\mathcal{P}=\{5,2,1\}$
    of $M$. Then $\mathcal{M}$ is matched to $\mathcal{P}$.
  • Consider the basis $\mathcal{M}=\{3,4,5\}$. Then
    $\mathcal{N}=f(\mathcal{M})=\{3,2,1\},$
    which is not a basis of $M$. Choose $a_{2}=4$ and $x=5$ as in Claim 3.3, and form the basis
    $\mathcal{P}=\{3,5,1\}$
    of $M$. Then $\mathcal{M}$ is matched to $\mathcal{P}$.
Since every basis $\mathcal{M}$ of $M$ is matched to a basis of $M$, it follows that $M$ is matched to itself.
Example 3.8.
In this example we give a non-paving matroid over an abelian group for which the construction used in the proof of Theorem 3.2 breaks down.
Let $G=\mathbb{Z}$ and let
$E=\{1,3,7,9\}.$
Define a matroid $M$ of rank $3$ on $E$ by declaring its bases to be exactly
$\mathcal{B}(M)=\bigl\{\{1,7,9\},\{3,7,9\}\bigr\}.$
It is immediate that this is a matroid: the collection $\mathcal{B}(M)$ is nonempty, and the basis exchange axiom is trivially satisfied since the only two bases differ by exchanging $1$ and $3$.
Since the $2$-subset $\{1,3\}$ is contained in no basis, it is dependent. Hence $M$ is not paving.
Now define a matching $f:E\to E$ in the group setting by
$f(1)=9,\qquad f(3)=1,\qquad f(7)=3,\qquad f(9)=7.$
Consider the basis
$\mathcal{M}=\{3,7,9\}.$
Following the construction in the proof of Theorem 3.2, set
$\mathcal{N}=f(\mathcal{M})=\{1,3,7\}.$
Then $\mathcal{N}$ is not a basis of $M$, since $\mathcal{N}\notin\mathcal{B}(M)$.
Now take $a_{3}=9$ and $x=9\in E\setminus\mathcal{N}$. Since
$9+9=18\notin E,$
the analogue of Claim 3.3 holds. Replacing the corresponding element $7$ of $\mathcal{N}$ by $x=9$, we obtain
$\mathcal{P}=\{1,3,9\}.$
Again, $\mathcal{P}$ is not a basis of $M$, since $\mathcal{P}\notin\mathcal{B}(M)$.
Therefore, the replacement step used in the proof of Theorem 3.2 may fail for non-paving matroids. This shows that the paving assumption is needed in order to conclude that the modified set is again a basis.